A boat traveled 210 miles downstream and back. The trip downstream took 10 hours. The trip back took 70 hours. What is the speed of the boat in still water? What is the speed of the current?
Answer by mananth(12270)

distance traveled downstream = 210 miles
Time taken = 10 hours
downstream speed = 210/10 = 21 mph
..
distance traveled upstream = 210 miles
Time taken = 70 hours
upstream speed = 210/70 = 3 mph

let speed of boat in still water be x mph
let speed of current be y mph
.
x+y=21
x-y=3
.
add the two equations
2x=24
x=12 speed of boat in still water
plug the value of x in any equation above
y=9 mph speed of current